# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

### v7

@@ Abstract
Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily. Efficiently managing, processing, and displaying this data requires specialized expertise and is often tedious and repetitive. Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution. However, LLMs are not adept at handling complex numerical computations and table manipulations and are also constrained by a limited context budget. Based on this, we propose Data-Copilot, a data analysis agent that autonomously performs querying, processing, and visualization of massive data tailored to diverse human requests. The advancements are twofold: First, it is a code-centric agent that receives human requests and generates code as an intermediary to handle massive data, which is quite flexible for large-scale data processing tasks. Second, Data-Copilot involves a data exploration phase in advance, which explores how to design more universal and error-free interfaces for real-time response. Specifically, it actively explores data sources, discovers numerous common requests, and abstracts them into many universal interfaces (code modules) for daily invocation. When deployed in real-time requests, Data-Copilot only needs to invoke these pre-designed interfaces, transforming raw data into visualized outputs (e.g., charts, tables, and text) that best match the user’s intent. Compared to generating code from scratch, invoking these pre-designed and compiler-validated interfaces can significantly reduce errors during real-time requests. Additionally, interface workflows are more efficient and offer greater interpretability than code. We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news, demonstrating promising application prospects. Quantitative evaluations indicate that, compared to direct code generation and other agent strategies, our exploration-deployment strategy addresses human requests more accurately and efficiently, i.e., with higher success rates and lower token consumption.
Figure 1: We present two strategies for using LLMs to automate data analysis. Upper: Directly using LLMs to read and process massive data is limited by context length, the LLM’s capabilities, and output format constraints. Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks. It actively explores how to design more universal and error-free interface modules, improving the success rate of real-time requests. Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow to fulfill human intent.
@@ 1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption. However, these data-related tasks often require specialized knowledge and tedious manual labor.
@@
Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022), especially when combined with techniques such (Wei et al., 2022b; Kojima et al., 2022) has demonstrated their capability to handle complex tasks. Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow that performs data analysis and outputs visualization results that best match user expectations?
@@
An intuitive solution is to directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023), just as we do with text. However, several challenges must be considered: (1) Due to the input context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text. Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources. (2) Data processing is very complex, involving many tedious numerical calculations and intricate table manipulations. LLMs are not adept at performing these tasks. (3) An LLM can only output text, whereas we usually prefer charts or tables when analyzing data. These challenges significantly constrain the application of LLMs in data-related tasks.
@@
Recently, many works have explored alternative solutions. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024) and DB-GPT (Xue et al., 2023) apply LLMs to data science domain like Text2SQL. Besides, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). Data Interpreter (Hong et al., 2024) proposes a code-based agent for data science through the plan-code-verify paradigm. These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent. Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and output multi-modal results (e.g., chart, table, text) for humans, reducing tedious human labor and eliminating the need for direct data access by the LLM. However, creating a reliable code agent for various data analysis tasks is also not an easy feat. Firstly, LLM-generated code often contains formatting errors, logical inconsistencies, and code hallucinations, such as fabricating non-existent functions, which may prevent it from being deployed in real-world environments. Secondly, most code-based agent generates code for solving user requests, without considering the relations between different requests and only addressing each issue independently.
@@
Based on this observation, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis. First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches the user’s intent. The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks. Besides, Data-Copilot also incorporates a data exploration phase. During this phase, it actively explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance. In a real-world deployment, Data-Copilot flexibly invokes pre-design interface modules to handle user requests, creating an interface workflow that transforms raw data into desired visualization results.
@@
This design can bring three advantages. Firstly, this exploratory process allows Data-Copilot to analyze connections between data, design general interfaces for similar requests, and pre-validate their correctness, thereby reducing errors in real-time responses. Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving generation efficiency. Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
As shown in Figure 1, Data-Copilot contains three key steps for data exploration:
  1. 1.
@@
    Explore data and synthesize requests: Data-Copilot discovers potential requests and a broader range of human need scenarios from data. It involves a "self-exploration" process to generate massive synthesized requests based on all data schemas and seed requests.
  2. 2.
    Interface design and test: Data-Copilot generates modular code for request solving. Each interface is a functional module, and test cases are generated to verify the designed interface.
  3. 3.
@@
    Interface optimization: To improve the versatility of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces based on the compiler’s feedback.
@@
After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different requests, Data-Copilot can flexibly deploy various invocation structures, such as step-by-step serial workflows, parallel, or loop. It can even output a hybrid form of interface workflows and raw code for some "unfamiliar" requests. Our contributions are threefold:
@@
Figure 2: Overview of Data-Copilot. Data Exploration: First, it performs a self-exploration process to uncover potential human requests from data sources. Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces one by one. This exploration-design-optimization process is operated in advance. Workflow Deployment: Upon receiving real-time requests, Data-Copilot can quickly invoke existing interfaces and deploy a workflow for familiar requests, or flexibly combine interfaces and raw code for "uncovered" requests.
  1. 1.
@@
    We propose a code-centric agent, Data-Copilot, for massive data analysis tasks. It leverages LLM and its code generation capabilities to automate data querying, processing, and visualization tailored to human requests, reducing tedious human labor.
  2. 2.
@@
    We decouple code generation into two phases: data exploration and workflow deployment. In the exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data. These interfaces enhance the success rate and efficiency of real-time responses. In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests.
  3. 3.
@@
    We open-source Data-Copilot for diverse Chinese financial data analysis, including stocks, funds, and live news. Quantitative evaluations indicate that our agent outperforms other agent strategies, with higher success rates and lower token consumption. Besides, interface workflows are more convenient for human inspection and interaction, offering stronger interpretability.
@@ 2 Data-Copilot
Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent. Data-Copilot operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-exploration process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these synthesized requests, it abstracts universal interface modules for each request (Step 2: design). After that, it optimizes similar interfaces and tests their correctness, ensuring each interface is correct and universal (Step 3: testing and optimization). The whole exploration-design-optimization process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
When faced with real-time requests (Section 2.2), Data-Copilot can flexibly invoke pre-designed interfaces or directly generate raw code based for the user’s request. In most cases, existing interfaces can cover the majority of real-world requests, significantly enhancing both success rate and response speed. We provide a detailed prompt for two phases in Section F.1 and F.2.
@@ 2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe what data is available, understand the data formats, and determine how to use them. Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability. Similar to this, Data-Copilot also autonomously explores data and derives insights from vast data sources, including the relations between the data, and the potential requests associated with the data. Then Data-Copilot abstracts these exploratory insights into numerous reusable code components (interfaces), testing their correctness and optimizing their generality. This process of exploring data, identifying common requests, designing general interfaces, and testing and optimizing their performance is conducted in advance on its own.
Self-Exploration Request
@@
To explore data and mine insights, we design a self-exploration process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs should generate requests specifically based on provided data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.

### v8

@@ Abstract
Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily. Efficiently managing, processing, and visualizing such data is labor-intensive and frequently necessitates specialized expertise. Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution. However, LLMs are not adept at handling complex numerical computations and table manipulations, and they are further constrained by a limited length context. To bridge this, we propose Data-Copilot, a data analysis agent that autonomously performs data querying, processing, and visualization tailored to diverse human requests. The advancements are twofold: First, it is a code-centric agent that leverages code as an intermediary to process and visualize massive data based on human requests, achieving automated large-scale data analysis. Second, Data-Copilot involves a data exploration phase in advance, which autonomously explores how to design universal and error-free interfaces from data, reducing the error rate in real-time responses. Specifically, It imitates common requests from data sources, abstracts them into universal interfaces (code modules), optimizes their functionality, and validates effectiveness. For real-time requests, Data-Copilot invokes these interfaces to address user intent. Compared to generating code from scratch, invoking these pre-designed and well-validated interfaces can significantly reduce errors during real-time requests. We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news. Quantitative evaluations indicate that our exploration-deployment strategy addresses human requests more accurate and efficiently, with good interpretability.
Data-Copilot: Bridging Billions of Data and Humans with
Autonomous Workflow
Wenqi Zhang1, Yongliang Shen1, Zeqi Tan1, Guiyang Hou1, Weiming Lu1, Yueting Zhuang1 1College of Computer Science and Technology, Zhejiang University {zhangwenqi, luwm}@zju.edu.cn Project Page: https://github.com/zwq2018/Data-Copilot
Figure 1: We compare two LLM strategies for automated data analysis. Upper: LLM’s capabilities, context length, and output format constraints limit the use of LLMs to process massive data. Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks. It explores how to design more universal and error-free interface modules, improving the success rate of real-time requests. Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow for human intent.
@@ 1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption. However, these data-related tasks often require tedious manual labor and specialized knowledge.
@@
Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022) and techniques (Wei et al., 2022b; Kojima et al., 2022) have demonstrated the capability to handle complex tasks. Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow performing data analysis and visualization that best matches user expectations?
@@
An intuitive solution is to treat data as a special text, i.e., directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023). However, as shown in Figure 1, several challenges must be considered: (1) Due to context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text. Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources. (2) Data processing is complex, involving many tedious numerical calculations and intricate table manipulations. LLMs are not adept at performing these tasks. (3) Data analysis typically requires visualizing the results of data processing, whereas LLMs are limited to generating text output. These challenges constrain the application of LLMs in data-related tasks.
@@
Recently, many agent-based designs have explored alternative solutions (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b; Nejjar et al., 2023; Li et al., 2024b). LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on exploring insight from data. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024b), DB-GPT (Xue et al., 2023) and TAG (Biswal et al., 2024) apply LLMs to data science domain like Text2SQL. Data Interpreter (Hong et al., 2024) proposes a plan-code-verify paradigm for automating machine learning tasks. These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent. Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and visualize its results (e.g., chart, table, text) for humans. However, creating a code agent that can be used in real-world data analysis tasks is far from an easy feat. ① LLMs struggle to generate high-quality, error-free code in a single attempt, often containing format errors, logical inconsistencies, or fabricating non-existent functions. ② Although the inference speed of LLMs has significantly improved, generating lengthy code still consumes a considerable amount of time and tokens. These two challenges—high code error rate and inefficient inference—must be addressed urgently.
@@
To address this, we observe most human requests are either similar or inherently related. By abstracting common demands into interfaces and validating their functionality in advance, we can significantly improve both the success rate and efficiency of real-time deployment. Therefore, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis. First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches user’s intent. The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks. Besides, Data-Copilot also incorporates a data exploration and interface design phase. It autonomously explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance. In real-world deployment, Data-Copilot flexibly invokes pre-design interface modules for most requests, deploying a well-verified interface workflow for data processing and visualization.
@@
Data-Copilot brings three advantages. Firstly, this exploratory process allows Data-Copilot to analyze and summarize the inherent connections between human requests, design general interfaces for similar requests, and pre-validate their correctness, reducing errors in real-time responses. Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving inference efficiency. Lastly, compared to lengthy code, these interfaces provide greater interpretability, since it is easier for human reading and interaction. To achieve this, we contains three steps (Figure 1) when self exploration:
@@
Explore data and Synthesize requests: Data-Copilot discovers potential requests and a broader range of human need from data. It involves a "self-exploration" process to generate massive requests based on all data schemas and seed requests.
Interface Design and Test: It designs modular interface from synthesized requests, with test cases automatically generated for verification.
@@
Interface Optimization: To improve versatility, it merges similar interfaces and also revises erroneous interfaces using compiler’s feedback.
@@
After exploration-design-optimization, Data-Copilot designs many general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different requests, Data-Copilot can flexibly deploy various invocation structures, such as step-by-step serial workflows, parallel, or loop. It can even output a hybrid form of interface workflows and raw code for these "unfamiliar" requests. Our contributions are threefold:
@@
Figure 2: Overview of Data-Copilot. Data Exploration: First, it performs a self-exploration process to uncover potential human requests from data sources. Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces. This exploration-design-optimization process is operated in advance. Workflow Deployment: Upon receiving real-time requests, Data-Copilot invokes existing interfaces and deploys a workflow for familiar requests, or flexibly combines interfaces and raw code for "uncovered" requests.
@@
  • We propose a code-centric agent, Data-Copilot, for automated data analysis and visualization. It leverages LLM’s code generation abilities for data querying, processing, and visualization, reducing tedious human labor.
@@
  • We decouple code generation into two phases: data exploration and workflow deployment. In exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data. In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests. This design enhances the success rate and efficiency of real-time responses.
@@
  • We open-source Data-Copilot for Chinese financial data analysis, including stocks, funds, and live news. Quantitative evaluations indicate our agent outperforms other strategies, with higher success rates and lower token consumption. Besides, interface workflows are more convenient for human inspection and interaction, offering interpretability.
@@ 2 Data-Copilot
Data-Copilot is a code-centric agent capable of performing data analysis and visualization based on human instructions. It operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-exploration process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these synthesized requests, it abstracts many universal interface modules (Step 2: design). After that, it optimizes similar interfaces and tests their correctness, ensuring each interface is correct and universal (Step 3: testing and optimization). The whole process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
When faced with real-time requests (Section 2.2), Data-Copilot can flexibly invoke pre-designed interfaces or directly generate raw code based for the user’s request. In most cases, existing interfaces can cover the majority of real-world requests, significantly enhancing both success rate and response speed. We provide a detailed prompt for two phases in Section E.3 and E.4.
@@ 2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe the available data, understand the data formats, and learn how to access them. Subsequently, humans often design generic modules to simplify the code logic and test these modules for usability. Similar to this, Data-Copilot also autonomously explores data and derives insights from vast data sources, including the relations between the data, and the potential requests associated with the data. Then Data-Copilot abstracts these exploratory insights into numerous reusable code components (interfaces), testing their correctness and optimizing their generality. This process of exploring data, identifying common requests, designing general interfaces, and testing and optimizing their performance is conducted in advance on its own.
@@
Self-Exploration Request. To explore data and mine insights, we design a self-exploration process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs should generate requests specifically based on provided data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.